# Prospects for High-Frequency Gravitational-Wave Detection with GEO600

### v2

@@ III.1.2 Detection Prospects
Our figure of merit to compare the detection capabilities of our detuned GEO600 configurations against LIGO is the sky-averaged SNR produced by a characteristic high-frequency source in each detector. Since the boson clouds produce a monochromatic signal, the sky-averaged SNR is given by [6] as
@@
[…]
Both the tuned GEO600 and aLIGO would be able to detect vector boson clouds over a large parameter space, though aLIGO generally tends to outperform GEO600 due to its superior broadband sensitivity. However, over a range of boson rest masses (and therefore characteristic GW frequencies), GEO600 can gain a competitive advantage over aLIGO in a narrow band around that characteristic frequency with a careful choice of the MSR detuning angle. Indeed, at frequencies above $\sim 6.5$ kHz, there exists some detuned configuration of GEO600 which can attain greater SNR for boson cloud GWs than aLIGO. Furthermore, for frequencies $\gtrsim 15$ kHz, GEO600 can obtain an SNR of eight with a week of integration time while aLIGO is not able to reach this threshold. By detuning the MSR and observing at that frequency range for a week, GEO600 can search for GWs sourced from boson clouds from $500$ Hz to $31.3$ kHz, with the latter frequency being obtained with a detuning angle of $\phi=45\degree$ (not plotted). Note that the pattern of high-frequency narrow bands in Fig. 5 is only shown up to $\phi=35\degree$, which results in a detection bandwidth centered at $24.5$ kHz.
@@ III.2.1 GW Strain Model
We simulate GWs from SSM compact binary mergers within the mass range of $10^{-6}~M_{\odot}$ to $1~M_{\odot}$. Note that the detection of GWs sourced from compact objects within this mass range would not be provide the sufficient parameters to distinguish that the object is a PBH specifically. In addition, if the source were a PBH, a detection within this mass range could only provide evidence that PBHs constitute some fraction of dark matter [76]. We compute the characteristic strain from the binary merger of SSM compact objects using the inspiral, merger, and ringdown waveforms available in PyCBC’s IMRPhenomD package [128]. For simplicity, we assume an inclination angle of $0\degree$, making $\tilde{h}_{+}(f)\!=\!\tilde{h}_{\times}(f)$ and $\tilde{h}(f)=\mathcal{R}(f)\tilde{h}_{0}(f)$ (see Appendix A.3). We calculated the sky-averaged SNR using the chirp formula given by [6] as
@@
Figure 6: Strain sensitivity $A_{h}(f)$ of the aLIGO design (light blue) and scanned GEO600 (light gray). Three different GW waveforms are plotted as $2|\tilde{h}_{+}(f)|\sqrt{f}$. The GW amplitudes are from compact binary mergers of three different sub-solar masses, observed from a distance of $30$ kPc.
@@ III.2.2 Detection Prospects
Our figure of merit to compare the detection capabilities of our detuned GEO600 configurations against LIGO is the sky-averaged SNR produced by a characteristic high-frequency SSM compact binary merger in each detector. In Fig. 6, we show the GW strain of compact mergers with individual BH masses of $10^{-1}$, $10^{-2}$, and $10^{-3}~M_{\odot}$, located a distance of 30 kPc from the detector. As the chirp mass decreases, $f_{\text{min}}$ increases while the GW strain is reduced. The frequency-domain signal is calculated with an observation time of one week, and the SNR is related to the area in the plot between the strain and noise spectral functions. We find that there are not SSM binaries which can be detected easier with a detuned GEO600 detector. Over the frequency range where the detuning-scanned GEO600 has a better sensitivity than aLIGO, the strain from the merger is weak, and thus the accumulated SNR is so little that it does not sway the total SNR in favor of GEO600. While the time domain signal is stronger at high frequencies, the early inspiral dominates the SNR because it has more cycles in band. If there were a SSM binary that were very slowly evolving in frequency at a timescale of $\sim$ months, then detuning GEO to the exact frequency could improve the sensitivity compared with LIGO. For nearly monochromatic SSM signals, one would see a similar improvement as to what we saw in Sec. III.1. Thus, GEO600 is not an optimal ground-based interferometer for the detection of most SSM compact binary mergers.
@@ IV Conclusion
In this work, we investigated how the high-frequency sensitivity of GEO600 can be scanned by detuning the location of the MSR. The MSR creates an optical cavity that enables narrow-band resonant amplification of GW-driven sideband fields. Detuning the MSR changes which frequencies satisfy the resonant condition, allowing GEO600 to scan over a range of high frequencies with broadband sensitivity. We estimated the SNR in the experiment for characteristic high-frequency GW sources located within the Milky Way. This setup of detuning GEO600 does not result in an improvement in searches for SSM binaries due to their long inspiral in low frequencies, and the bandwidth being $\sim 2$ kHz. However, GEO600 could outperform the aLIGO design sensitivity for the detection of GWs sourced from nearly monochromatic sources like boson clouds within the galaxy. The interferometer would be able to detect GWs sourced from vector boson clouds at frequencies between $500$ Hz to $31.3$ kHz, gaining the advantage over the aLIGO design sensitivity at $15.1$ kHz and over O4 LIGO at $10$ kHz.
@@ A.2 Folded Michelson Interferometer
where the extra factor of 2 comes from the DC antenna pattern normalization. In this equation, we stress that $F_{A}^{\mathrm{mich}}(\hat{n},f)$ is a Michelson interferometer of length 600m while $F_{A}^{\mathrm{fold}}(\hat{n},f)$ is a folded Michelson interferometer of length 600m+600m. So while it is understood that GEO600 has an effective arm length of 1200m because of the folding, the high-frequency antenna patterns of GEO600 behave more like a 600m detector times a factor. One notable feature of the folded detector is that there is zero response for every angle at frequency values of
@@ B.6.3 Test Mass Coating Thermal Noise
Figure 8: Strain sensitivity $A_{h}(f)$ of GEO600 with a tuned MSR (left) and an anti-tuned MSR (right), including all modeled noise sources and the total sensitivity, both with and without sky-averaging (though individual sources are only shown for normal incidence). BM: Brownian motion, TE: thermo-elastic, TR: thermo-refractive.

### v3

@@ III.1.2 Detection Prospects
Our figure of merit to compare the detection capabilities of our detuned GEO600 configurations against LIGO is the sky-averaged SNR produced by a characteristic high-frequency source in each detector. Since the boson clouds produce a monochromatic signal, the sky-averaged SNR is given by [99]
@@
In Fig. 5, we show contours of SNR equal to 8 for vector and scalar boson clouds for different interferometer configurations, representing regions of this parameter space which are detectable by each system. The plot shows the feasibility of certain configurations of GEO600 detecting GWs for different combinations of boson rest energies and BH mass. We also present a horizontal cross-section of the contour plot, which shows the SNR of the interferometer at different boson rest energies for a given BH mass $m_{\text{BH}}=0.3M_{\odot}$.
Both the tuned GEO600 and aLIGO would be able to detect vector boson clouds over a large parameter space, though aLIGO generally tends to outperform GEO600 due to its superior broadband sensitivity. However, over a range of boson rest masses (and therefore characteristic GW frequencies), GEO600 can gain a competitive advantage over aLIGO in a narrow band around that characteristic frequency with a careful choice of the MSR detuning angle. Indeed, at frequencies above $\sim 6.5$ kHz, there exists some detuned configuration of GEO600 which can attain greater SNR for boson cloud GWs than aLIGO.
Furthermore, for frequencies $\gtrsim 15$ kHz, GEO600 obtains an SNR of eight with a week of integration time while aLIGO is not able to reach this threshold. By detuning the MSR and observing at that frequency range for a week, GEO600 can search for GWs sourced from boson clouds from $500$ Hz to $31.3$ kHz, with the latter frequency being obtained with a detuning angle of $\phi=45\degree$. Note that the pattern of high-frequency narrow bands in Fig. 5 is only shown up to $\phi=35\degree$, which results in a detection bandwidth centered at $24.5$ kHz.
@@ III.2.1 GW Strain Model
We simulate GWs from SSM compact binary mergers within the mass range of $10^{-6}~M_{\odot}$ to $1~M_{\odot}$. Note that the detection of GWs sourced from compact objects within this mass range would not be provide the sufficient parameters to distinguish that the object is a PBH specifically. In addition, if the source were a PBH, a detection within this mass range could only provide evidence that PBHs constitute some fraction of dark matter [61]. We compute the characteristic strain from the binary merger of SSM compact objects using the inspiral, merger, and ringdown waveforms available in PyCBC’s IMRPhenomD package [39]. For simplicity, we assume an inclination angle of $0\degree$, making $\tilde{h}_{+}(f)\!=\!\tilde{h}_{\times}(f)$ and $\tilde{h}(f)=\mathcal{R}(f)\tilde{h}_{0}(f)$ (see Appendix A.3). We calculated the sky-averaged SNR using the chirp formula given by [99]
@@
Figure 6: Strain sensitivity $A_{h}(f)$ of the aLIGO design (light blue) and scanned GEO600 (light gray). Three different GW waveforms are plotted as $2|\tilde{h}_{+}(f)|\sqrt{f}$. The GW amplitudes are from compact binary mergers of three different sub-solar masses, observed from a distance of $30$ kPc. Note that all strain sensitivities assume a GW sky location of $\theta_{\hat{n}}=15$ and $\phi_{\hat{n}}=0$.
@@ III.2.2 Detection Prospects
Our figure of merit to compare the detection capabilities of our detuned GEO600 configurations against LIGO is the sky-averaged SNR produced by a characteristic high-frequency SSM compact binary merger in each detector. In Fig. 6, we show the GW strain of compact mergers with individual BH masses of $10^{-1}$, $10^{-2}$, and $10^{-3}~M_{\odot}$, located a distance of 30 kPc from the detector. As the chirp mass decreases, $f_{\text{min}}$ increases while the GW strain is reduced. The frequency-domain signal is calculated with an observation time of one week, and the SNR is related to the area in the plot between the GW waveform $\tilde{h}(f)$ and the strain sensitivity corresponding to a single interferometer configuration $A_{h}(f)$. Note that FIG. 6 only shows the scanned sensitivity of GEO600, which is the ASD optimized pointwise in frequency space over MSR detuning angles at all frequencies and not a realizable configuration itself.
We find that there are not SSM binaries which can be detected easier with a detuned GEO600 detector. Over the frequency range where the detuning-scanned GEO600 has a better sensitivity than aLIGO, the strain from the merger is weak, and thus the accumulated SNR from any single detuned configuration is so little that it does not sway the total SNR in favor of GEO600. While the time domain signal is stronger at high frequencies, the early inspiral dominates the SNR because it has more cycles in band. If there were a SSM binary that were very slowly evolving in frequency at a timescale of $\sim$ months, then detuning GEO to the exact frequency could improve the sensitivity compared with LIGO. For nearly monochromatic SSM signals, one would see a similar improvement as to what we saw in Sec. III.1. Thus, GEO600 is not an optimal ground-based interferometer for the detection of most SSM compact binary mergers.
@@ IV Conclusion
In this work, we investigated how the high-frequency sensitivity of GEO600 can be scanned by detuning the location of the MSR. The MSR creates an optical cavity that enables narrow-band resonant amplification of GW-driven sideband fields. Detuning the MSR changes which frequencies satisfy the resonant condition, allowing GEO600 to scan a resonant feature with a bandwidth of $\sim 2$ kHz over a range of high frequencies. We estimated the SNR in the experiment for characteristic high-frequency GW sources located within the Milky Way. This setup of detuning GEO600 does not result in an improvement in searches for SSM binaries due to their long inspiral in low frequencies, and the strain from the merger being weak at high frequencies. However, GEO600 could outperform the aLIGO design sensitivity for the detection of GWs sourced from nearly monochromatic sources like boson clouds within the galaxy. By specifying a MSR detuning angle between $0\degree$ and $45\degree$, GEO600 would be able to detect GWs sourced from vector boson clouds at any $\sim 2$ kHz band of frequencies between $500$ Hz to $31.3$ kHz, gaining the advantage over the aLIGO design sensitivity at $15.1$ kHz and over O4 LIGO at $10$ kHz.
@@ A.2 Folded Michelson Interferometer
where the extra factor of 2 comes from the DC antenna pattern normalization. In this equation, we stress that $F_{A}^{\mathrm{mich}}(\hat{n},f)$ is a Michelson interferometer of length 600600m while $F_{A}^{\mathrm{fold}}(\hat{n},f)$ is a folded Michelson interferometer of length 600600m ++ 600600m =1200=1200m. So while it is understood that GEO600 has an effective arm length of 1200m because of the folding, the high-frequency antenna patterns of GEO600 behave more like a 600m detector times a factor. One notable feature of the folded detector is that there is zero response for every angle at frequency values of
@@ B.6.3 Test Mass Coating Thermal Noise
Figure 8: Strain sensitivity $A_{h}(f)$ of GEO600 with a tuned MSR (left) and an anti-tuned MSR (right), including all modeled noise sources and the total sensitivity, both with and without sky-averaging. Note that all individual sources, as well as the total sensitivity without sky-averaging, assume a normal-incidence GW sky location. BM: Brownian motion, TE: thermo-elastic, TR: thermo-refractive.